# Revealing the Origin of Neutrino Masses through Displaced Shower Searches in the CMS Muon System

### v1

@@ A.3 Neutrino Sector
${\cal M}=\begin{pmatrix}0&m_{D}\\ m^{T}_{D}&M_{N}\end{pmatrix},$ (A.9)
Here, $v=\langle H^{0}\rangle\approx 246$ GeV and $v_{B-L}=\langle\Phi\rangle$ are the vacuum expectation values for electroweak and $B-L$ symmetry breaking, respectively. In the seesaw limit, $M_{N}\gg m_{D}$, the light neutrino masses are
$m_{\nu}\sim-m_{D}M^{-1}_{N}m^{T}_{D}.$ (A.10)
The flavour and mass eigenstates of the light and heavy neutrinos are related as
@@
The mixing and the light neutrino masses are constrained by oscillation experiments, namely, the charged current lepton mixing $U\approx U_{\text{PMNS}}$, apart from small non-unitarity corrections. For simplicity, we assume the presence of a single RHN $N$ mixing with a single lepton flavour at a time. This relates the light neutrino mass scale $m_{\nu}$ with the RHN mass $m_{N}$ as $m_{\nu}=|V_{\ell N}|^{2}m_{N}$. While this simplification does not allow to describe the full light neutrino phenomenology, the mass should not exceed limits on the absolute neutrino mass scale currently set by the KATRIN experiment as $m_{\nu}<0.45$ eV at 90 % CL Aker et al. (2024). While not a strict lower limit, observations of the solar neutrino oscillation length indicate a smallest non-vanishing neutrino mass scale of $\sqrt{\Delta m^{2}_{\text{sol}}}=9\times 10^{-3}$ eV Esteban et al. (2020). We thus take the range $0.459\times 10^{-3}~{}\text{eV}<|V_{\ell N}|^{2}m_{N}<$ 0.45 eV as a target to probe the canonical seesaw floor of neutrino mass generation.
@@
RHNs can be searched for through their mixing $V_{\ell N}$ with the active neutrinos, via their resulting participation in SM neutral and charged-current interactions, irrespective of the presence of a $B-L$ gauge interaction. This leads to a wide range of constraints at colliders, in beam dump experiments, meson decay searches, etc.. In the figures in the main text, we display current constraints compiled in Bolton et al. (2020) for comparison with our production mechanisms. The focus of future efforts is on the RHN lifetime frontier and we display the projected sensitivities of SHiP Ahdida et al. (2019), CMS Drewes and Hajer (2020) and FCC-ee Blondel et al. (2022) as representative examples in our parameter space of interest.

### v2

@@ A.3 Neutrino Sector
${\cal M}=\begin{pmatrix}0&m_{D}\\ m^{T}_{D}&M_{N}\end{pmatrix}.$ (A.9)
In the seesaw limit, $M_{N}\gg m_{D}$, the light neutrino masses are
$m_{\nu}=-m_{D}M^{-1}_{N}m^{T}_{D},$ (A.10)
and the flavour and mass eigenstates of the light and heavy neutrinos are related as
@@
The mixing and the light neutrino masses are constrained by oscillation experiments, namely, the charged current lepton mixing $U\approx U_{\text{PMNS}}$, apart from small non-unitarity corrections. For simplicity, we assume the presence of a single RHN $N$ mixing with a single lepton flavour at a time. This relates the light neutrino mass scale $m_{\nu}$ with the RHN mass $m_{N}$ as $m_{\nu}=|V_{\ell N}|^{2}m_{N}$. While this simplification does not allow describing the full light neutrino phenomenology, the mass should not exceed limits on the absolute neutrino mass scale currently set by the KATRIN experiment as $m_{\nu}<0.45$ eV at 90 % CL [6]. While not a strict lower limit, observations of the solar neutrino oscillation length indicate a smallest non-vanishing neutrino mass scale of $\sqrt{\Delta m^{2}_{\text{sol}}}=9\times 10^{-3}$ eV [7]. We thus take the range $0.459\times 10^{-3}~{}\text{eV}<|V_{\ell N}|^{2}m_{N}<$ 0.45 eV as a target to probe the canonical seesaw floor of neutrino mass generation.
@@
RHNs can be searched for through their mixing $V_{\ell N}$ with the active neutrinos, via their resulting participation in SM neutral and charged-current interactions, irrespective of the presence of a $B-L$ gauge interaction. This leads to a wide range of constraints at colliders, in beam dump experiments, meson decay searches, etc.. In the figures in the main text and below, we display current constraints compiled in [12] for comparison with our production mechanisms. The focus of future efforts is on the RHN lifetime frontier and we display the projected sensitivities of SHiP [126], CMS [127] and FCC-ee [128] as representative examples in our parameter space of interest.
A.4 Sensitivity of the CMS Displaced Shower Search
Figure 7: Sensitivity of $pp\to h\to NN$ (left) and $pp\to\Phi\to NN$ ($m_{\Phi}=150$ GeV, right) at 95% CL on the RHN mass $m_{N}$ and the active-sterile mixing $|V_{\ell N}|^{2}$ ($\ell=e,\tau$). The scenarios are for reinterpreting CMS data (red region), as well as at the 14 TeV HL-LHC with 3000 fb-1 using the CMS (blue contour) and soft (green contour) trigger strategies. The Higgs mixing is $\sin\alpha=0.24$ (red region and solid contours) and $0.14$ (dashed), and the $B-L$ Higgs VEV is $\langle\Phi\rangle=3.75$ TeV. The region labelled ‘Current Limits’ is excluded by sterile neutrino searches ($\ell=e$) while the contours labelled SHiP, CMS and FCC-ee give the projected sensitivity of planned experiments [12, 104]. The band labelled ‘Seesaw’ indicates $0.459\times 10^{-3}~{}\text{eV}<|V_{\ell N}|^{2}m_{N}<$ 0.45 eV.
As discussed in the main text, the CMS search for displaced showers in the muon endcap can be interpreted in terms of the $B-L$ gauge model via the processes $pp\to h\to NN$ and $pp\to\Phi\to NN$ with RHNs decaying in the muon system. For completeness, we show the resulting sensitivity in Fig. 7, with the left panel being identical to Fig. 2 in the main text.
Appendix B Interpretation in the RHN-extended SM Effective Field Theory
Instead of choosing a specific ultraviolet-complete scenario such as the $U(1)_{B-L}$ gauge model described above, the effects of heavy New Physics at SM scales and below can be categorized and interpreted within an Effective Field Theory (EFT) approach. Assuming that the SM particle content is extended by a single, sterile and relatively light RHN ($m_{N}\lesssim\Lambda_{\text{EW}}$) with no other exotic states present, the so-called RHN-extended SM EFT ($\nu_{R}$) [46, 47, 48, 49, 50] applies, and the Lagrangian can be expressed as
[…]
At the renormalizable level, it adds a kinetic term and a Majorana mass term for the RHN $\nu_{R}$ field as well as the usual Yukawa interactions $\lambda_{D}^{\alpha}$ to the SM lepton doublets. This gives rise to the usual seesaw mechanism, where we consider the single RHN $N\approx\nu_{R}$ with mass $m_{N}$ to generate a light neutrino mass scale $m_{\nu}=|V_{\ell N}|^{2}m_{N}$ as described in Sec. A.3. Low energy effects of the heavy New Physics are then captured by effective operators $\mathcal{O}_{n}$ of increasing mass dimension $n$, constructed from SM fields and the RHN. They are accompanied by dimensionless Wilson coefficients ${\cal C}_{n}$ quantifying the individual operator strengths, usually expected to be $|{\cal C}_{n}|=\mathcal{O}(1)$ unless additional symmetry considerations lead to a suppression. The set of operators includes those of the SMEFT (without RHN) and additional operators incorporating the RHN.
At the lowest order, $n=5$, there are two operators  [^2] There exists a third dimension-5 operator, ${\cal O}_{NB}=(\bar{\nu}_{R}^{c}\sigma^{\mu\nu}\nu_{R})B_{\mu\nu}=0$ that is usually considered but is absent for a single RHN, reflecting the fact that a Majorana fermion does not have a magnetic dipole moment.,
${\cal O}_{W}^{\alpha\beta}$ $=(\bar{L}^{\alpha,c}\cdot\tilde{H}^{*})(\tilde{H}^{\dagger}\cdot L^{\beta}),$
${\cal O}_{NH}$ $=(\bar{\nu}_{R}^{c}\nu_{R})(H^{\dagger}\cdot H).$ (A.13)
The first, ${\cal O}_{W}^{\alpha\beta}$, is the well-known Weinberg operator inducing light neutrino masses after EW symmetry breaking. In the $\nu_{R}$, it describes an additional contribution $\delta m_{\nu}^{\alpha\beta}=-\mathcal{C}_{W}^{\alpha\beta}v^{2}/\Lambda_{\text{NP}}$, beyond that from the RHN-induced seesaw. This could, e.g., arise from additional RHN states heavier than the EW scale, or from other neutrino mass generation mechanisms in the ultraviolet. The second operator, ${\cal O}_{NH}$, is the RHN equivalent of the Weinberg operator. It induces an additional RHN Majorana mass contribution after EW symmetry breaking, $\delta m_{N}=-\mathcal{C}_{NH}v^{2}/\Lambda_{\text{NP}}$. More importantly in our context, it also generates the $hNN$ vertex with coupling strength $\mathcal{C}_{NH}v/\Lambda_{\text{NP}}$.
The operator ${\cal O}_{NH}$ will thus induce the Higgs branching ratio
$\text{Br}(h\to NN)=\frac{1}{4\pi}\frac{|\mathcal{C}_{NH}|^{2}v^{2}}{\Lambda^{2}_{\text{NP}}}\frac{m_{h}}{\Gamma_{h}^{\text{SM}}}\left(1-\frac{4m_{N}^{2}}{m_{h}^{2}}\right)^{3/2},$ (A.14)
which is the equivalent of Eq. (4) in the main text. With the RHN decaying via the active-sterile mixing $V_{\ell N}$, as in the $U(1)_{B-L}$ model, the constraints from the CMS search and future sensitivities can thus be reinterpreted in terms of the operator scale $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|$ in the $\nu_{R}$. For this, we assume that other operators do not contribute significantly to the RHN decay width.
Figure 8: Lower limit on the $\nu_{R}$ operator scale $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|$ as a function of the RHN mass $m_{N}$ and the active-sterile mixing strength $|V_{\ell N}|^{2}$ using existing CMS data (top) as well as at the 14 TeV HL-LHC with 3000 fb-1 using the CMS (bottom left) and soft trigger strategies (bottom right). The other elements in the plots are as described in Fig. 7.
For a given RHN mass $m_{N}$ and active-sterile mixing strength $|V_{\ell N}|^{2}$ we determine the smallest operator scale $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|$ that yields no significant signal in the CMS muon system at 95% CL by reinterpreting it in terms of the Higgs branching ratio in Eq. (A.14). This is shown in Fig. 8, using the existing CMS data (top, identical to Fig. 3 in the main text) as well as at the 14 TeV HL-LHC with 3000 fb-1 using the CMS (bottom left) and soft trigger strategies (bottom right). As in the $U(1)_{B-L}$ scenario, the search is sensitive as long as the RHN is sufficiently long-lived, $0.5~{}\text{m}\lesssim L_{N}\lesssim 6$ m, to be detected in the CMS muon system. Within this band, there is a soft dependence on the RHN mass, with the sensitivity decreasing with larger $m_{N}$ until nearing the threshold $m_{N}=m_{h}/2$.
Figure 9: Best upper limit and projected sensitivity (maximized over $|V_{\ell N}|^{2}$) on the $\nu_{R}$ operator scale $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|$ as a function of the RHN mass $m_{N}$ from the existing CMS search (red), at the HL-LHC with the CMS trigger (blue) and at the HL-LHC with the soft trigger (green). The current limit on the operator scale from the Higgs signal strength [109] is shown as well. A contribution to the RHN mass from the operator $\mathcal{O}_{NH}$ at the level $|\delta m_{N}|=0.1m_{N}$ and $0.01m_{N}$ is indicated by the dashed curves.
In Fig. 9, we likewise show the best constraint on $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|$ for a given RHN mass $m_{N}$, comparing the existing CMS search with the future sensitivities at the HL-LHC with the CMS trigger and the soft trigger as described in the main text. The sensitivity peaks at $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|\gtrsim 300,500,3000$ TeV with CMS data, at the HL-LHC using the CMS trigger and at the HL-LHC using the soft trigger, respectively. We also show the limit from Higgs signal strength measurements, i.e., interpreting $h\to NN$ as an invisible decay, at $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|\gtrsim 33$ TeV [108, 109]. The sensitivity of the existing CMS displaced shower search surpasses this limit by an order of magnitude. This is expected to improve by a hundred-fold at the HL-LHC using the soft trigger, which has already been installed since the start of Run-3 [103]. As mentioned, the operator ${\cal O}_{NH}$ induces a correction to the bare mass of the RHN after EW symmetry breaking, $\delta m_{N}=-\mathcal{C}_{NH}v^{2}/\Lambda_{\text{NP}}$. For consistency, it should be small compared to the bare mass and corrections at the 10% and 1% level are indicated by the dashed curves in Fig. 9.
As the main difference to the interpretation in the ultraviolet-complete $U(1)_{B-L}$ gauge model, the sensitivity to $\Lambda_{NP}$ does not decrease for small $m_{N}$. This is because the $hNN$ coupling in the $U(1)_{B-L}$ model arises from the breaking of lepton number, generating the RHN Majorana mass, and thus the $hNN$ vertex is proportional to $m_{N}$. This is expected to be a generic feature if the source of lepton number violation can be directly connected to the RHN mass. It can be incorporated within the $\nu_{R}$ via the inclusion of a spurion [130], associated with lepton number, to reflect the additional symmetry and its controlled breaking in the ultraviolet sector. As a consequence, effective operators violating lepton number vanish if lepton number is conserved in the renormalizable Lagrangian, i.e., as $m_{N}\to 0$. The relevant Wilson coefficients are thus suppressed as $\mathcal{C}_{\text{LNV}}=\mathcal{C}^{\prime}_{\text{LNV}}m_{N}/\Lambda_{\text{NP}}$ with $|\mathcal{C}^{\prime}_{\text{LNV}}|=\mathcal{O}(1)$. The $h\to NN$ branching ratio then scales as
$\text{Br}(h\to NN)=\frac{1}{4\pi}\frac{|\mathcal{C}^{\prime}_{NH}|^{2}m_{N}^{2}v^{2}}{\Lambda^{4}_{\text{NP}}}\frac{m_{h}}{\Gamma_{h}^{\text{SM}}}\left(1-\frac{4m_{N}^{2}}{m_{h}^{2}}\right)^{3/2}.$ (A.15)
In this context, we can further reinterpret the sensitivity to $\Lambda_{\text{NP}}$ by making the substitution $|\mathcal{C}_{NH}|/\Lambda_{\text{NP}}\to|\mathcal{C}^{\prime}_{NH}|m_{N}/\Lambda^{2}_{\text{NP}}$. The previous sensitivities $\Lambda_{\text{NP}}/|\mathcal{C}_{NH}|=300,500,3000$ TeV for the CMS data, the HL-LHC (CMS trigger) and the HL-LHC (soft trigger) are then mapped to $\Lambda^{\prime}_{\text{NP}}/|\mathcal{C}_{NH}|=3.5,4.5,11$ TeV, respectively, for $m_{N}=40$ GeV. This is compatible with the sensitivity with respect to the $B-L$-breaking scale and the $Z^{\prime}$ mass in the $U(1)_{B-L}$ gauge model discussed in Sec. A.4.
Other $\nu_{R}$ operators at higher dimensions [50] may also contribute to the RHN and Higgs decay, and could be probed similarly. At dimension-6, operators coupling the RHN with the Higgs are [109]
${\cal O}_{LNH}^{\alpha}=(\bar{L}_{\alpha}\cdot\tilde{H}\nu_{R})(H^{\dagger}\cdot H),\quad{\cal O}_{NHH}=(\bar{\nu}_{R}\gamma^{\mu}\nu_{R})(H^{\dagger}i\overleftrightarrow{D}_{\mu}H),$ (A.16)
and
${\cal O}_{N^{2}H^{4}}=(\bar{\nu}_{R}\nu_{R})(H^{\dagger}\cdot H)^{2},$
${\cal O}_{ND1}=(\bar{\nu}_{R}D_{\mu}\nu_{R})(H^{\dagger}\overleftrightarrow{D}^{\mu}H),\quad{\cal O}_{ND2}=(\bar{\nu}_{R}\nu_{R})\big{(}(D_{\mu}H)^{\dagger}\cdot(D^{\mu}H)\big{)},$ (A.17)
at dimension-7 [50]. Here, $D_{\mu}$ is the SM covariant derivative. Displaced shower searches can also be sensitive to $h\to N\nu$ which, in addition to the usual active-sterile neutrino mixing contribution, involves further operators.